# Scene-Adaptive Motion Planning with Explicit Mixture of Experts and Interaction-Oriented Optimization

### v1

@@ Abstract
Despite over a decade of development, autonomous driving trajectory planning in complex urban environments continues to encounter significant challenges. These challenges include the difficulty in accommodating the multi-modal nature of trajectories, the limitations of single expert in managing diverse scenarios, and insufficient consideration of environmental interactions. To address these issues, this paper introduces the EMoE-Planner, which incorporates three innovative approaches. Firstly, the Explicit MoE (Mixture of Experts) dynamically selects specialized experts based on scenario-specific information through a shared scene router. Secondly, the planner utilizes scene-specific queries to provide multi-modal priors, directing the model’s focus towards relevant target areas. Lastly, it enhances the prediction model and loss calculation by considering the interactions between the ego vehicle and other agents, thereby significantly boosting planning performance. Comparative experiments were conducted using the Nuplan dataset against the state-of-the-art methods. The simulation results demonstrate that our model consistently outperforms SOTA models across nearly all test scenarios.
1 Introduction
@@
  • Propose a novel scene-specific query to enable the model to focus on area that the goal point is located. The query simultaneously provides prior information to the model at both the decision level and the modality level, allowing the model to achieve better training efficiency and performance with only a small number of queries.
@@ 3.1 Overview
Figure 1 illustrates the framework of the proposed model. The model receives the ego vehicle state information $F_{e}\in R^{D_{e}}$, dynamic object information $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, static object information $F_{s}\in R^{N_{s}\times D_{s}}$, and map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$. These inputs are processed through respective encoding modules to obtain D-dimensional features, specifically, [$E_{ego}\in R^{D},E_{agent}\in R^{N_{a}\times D},E_{static}\in R^{N_{s}\times D},E_{map}\in R^{N_{m}\times D}$]. The details of these input encoding modules will be elaborated in Section 3.2. It is worth noting that the positional and heading angle information in the input data are relative values defined in the ego vehicle’s coordinate system.
@@
Figure 1: Framework of EMoE-Planner
3.2 Feature Encoders
@@
As Equation 3 shows, SDE[7] and multi-layer perceptrons (MLP) are employed to encode $F_{e}$ and $F_{s}$. The ego status $F_{e}\in R^{D_{e}}$, where $D_{e}$ includes positional and heading information, is embedded into dimention $D$. Similarly, static object information $S\in R^{N_{s}\times D_{s}}$, where $N_{s}$ is the number of static objects and $D_{s}$ includes attributes such as position, heading, type, and shape, is also embedded into D for concatenation in later stages.
@@
Unlike EE and ES, the encoding of agent and map information requires consideration of both temporal and spatial compression. Equation 4 shows the process of EA and EM. Raw input of agent is $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, where $N_{a}$ is the number of agents, $T_{h}$ the historical time steps, and $D_{a}$ encompasses attributes such as position, heading, velocity, shape, and type. Dimention $D_{a}$ is first embedded into dimention $D$ with Fourier embeddings (FoPE)[21], followed by a compression of $N_{a}$ to a singular dimension via MLP-Mixer[43]. Fope effectively capture periodic patterns in the data, facilitating effective temporal compression. Subsequently, the MLP-Mixer employs dual layers of MLPs to perform interactions both within and across the temporal dimension. Similarly, the map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$ is also processed with two steps. But the first step is not with fope but a normal linear layer, which is better to handle spacial information.
$E_{agent}=\text{MLP-Mixer}(\text{FoPE}(F_{a})),E_{map}=\text{MLP-Mixer}(\text{Linear}(F_{m}))$ (4)
3.3 Explicit MOE
@@
Our approach deviates from traditional MoE models by introducing an architecture designed specifically for our task. Figure 2(a) depicts a standard MoE model where at each layer, a router selects top $N_{E}$ best-suited expert from all experts based on the input information, with $N_{E}=1$ in our case, to generate the output. However, this architecture has several limitations: the router lacks explicit constraints, making its selection logic difficult to interpret and prone to confusion in similar scenarios. It also risks an uneven distribution of expert utilization.
@@
Figure 2: Differences between normal MoE and EMoE
@@
3.4 Scene-specific query
@@
3.5 Interaction-oriented loss and prediction
3.5.1 Interaction-oriented loss
@@
The method for extracting interaction intervals is as follows: traverse each time step $t$ and assess all trajectory points of other agents for potential collisions based on their shapes and orientation angles. Subsequently, using the results from the collision checks, we determine the beginning and ending of interaction intervals, $t_{in},t_{out}\in[1:T_{f}]$. It is essential to clarify that the collisions mentioned herein do not pertain to actual physical interactions between vehicles but rather to instances where two vehicles occupy the same spatial location at different time steps. Furthermore, interaction labels (overtake/yield) are assigned based on the temporal sequence in which the vehicles arrive at the location. Specifically, if the ego vehicle at time $t_{e}$ “collides” with the i-th vehicle at time $t_{a}$, and $t_{e}<t_{a}$ (i.e., the ego vehicle passes through first), then $t_{in}=t_{e}$, and the interacting vehicle is labeled as "overtake".
For the i-th agent, the process of $t_{\text{in}}^{i}$ and $t_{\text{out}}^{i}$ is shown in Equation 7-8. The interaction interval is defined as the union of all ($t_{\text{in}}^{i}$,$t_{\text{out}}^{i}$) intervals, as shown in Equation 9.
$t_{\text{in}}^{i}=\min\left\{t_{a,i}\mid\exists t_{e}\in\{1,2,\ldots,T_{f}\},\|y_{a,i}^{t_{a,i}}-y_{e}^{t_{e}}\|<d_{\text{thr}}\right\}$ (7)
$t_{\text{out}}^{i}=\max\left\{t_{a,i}\mid\exists t_{e}\in\{1,2,\ldots,T_{f}\},\|y_{a,i}^{t_{a,i}}-y_{e}^{t_{e}}\|<d_{\text{thr}}\right\}$ (8)
@@
Based on the temporal-weighted regression loss in Equation refeq:reg, we set the weights of the ego trajectory points within the interaction intervals to 1 to highlight their importance.
@@
3.5.2 Interaction-oriented prediciton
@@
4 Experiments
4.1 Experimental Setup
@@
We trained all models on eight Nvidia H800 graphics cards, each with 80GB of VRAM, setting the batch size to 32 and training for 35 epochs. The training time for one million data is approximately 16 hours. The parameters used for training are listed in Table 1.
Table 1: Parameters Used in EMOE-Planner
Name Description Values Name Description Values
lr learning rate 1e-3 $T_{f}$ Future Steps 80
$L_{e}$ Num. scene encoder layers 4 $T_{h}$ History Steps 20
$L_{d}$ Num. decoder layers 4 $N_{E}$ Num. experts 7
$K_{a}$ Num. anchors 24 $k_{R}$ Weight decay factor 0.2
$D$ Dimention of features 128 $D_{h}$ Hidden dim of expert 512
4.2 Results
We assess the performance of all baseline algorithms and EMOE-Planner in NR simulations across Val14, Test14, and Test14-hard. Table 2 presents the evaluation outcomes while Table 3 details the metric results for the top two baseline methods and our method.
Table 2: Closed-loop Results. Higher scores indicate better performance.
@@ 4.2 Results
Table 3: Metrics Results.
@@
According to the results in Table 2, our model outperforms all other learning-based methods across all testing scenarios, particularly excelling in the val14 and test14-hard scenarios where it exceeds the state-of-the-art (SOTA) results by a minimum of 3 points. In the val14 scenarios, the performance of the EMOE-Planner closely approaches that of expert scores.Among rule-based methods, when implementing the same post-processing technique used in Pluto and DiffusionPlanner with our EMOE-Planner, it achieves the highest performance in the val14 and test14-hard scenarios, albeit slightly lower than DiffusionPlanner in the test14 scenario. It is noteworthy that the performance of EMOE-Planner with and without post-processing shows slight differences, suggesting that the trajectories produced by the model are of high quality, potentially diminishing the necessity for post-processing, aligning with our objectives.
@@
As indicated in Table 3, our model generally outperforms both Pluto and DiffusionPlanner in all metrics except speed. Notably, it excels in comfort and progress metrics. The low progress score of Pluto aligns with observations from simulations where Pluto’s trajectories tend to be conservative, often slower than expert trajectories. In the appendix, we provide detailed trajectories for various scenarios to give a comprehensive understanding of the situations discussed here. Our loss function accounts for vehicle interactions, which effectively optimizes the trajectory progress, particularly in overtaking scenarios, thus avoiding excessive conservatism while preventing collisions.
@@
In the EMOE architecture, the distribution of scenario quantities in the training data is crucial for model performance. Table 4’s second row details the scenario counts in the training dataset. During training, we capped each scenario type at 200,000 from a pool of 3,000,000 samples as discussed in section 4.1. The first four scenario types reached this cap, while the last three did not due to insufficient data in the 3M training database, resulting in a total of 1,083,804 samples used for training. We categorized the scenarios in val14 using the method outlined in the appendix and conducted separate simulation assessments for each scenario type, with results presented in Table 4.
Table 4: Results of different scenarios. ST-J, LT-J, RT-J means go straight, turn left and turn right in junctions. ST means go straight in non-junction areas. UT is u-turn and RA is roundabout.
Scenario Type ST-J LT-J RT-J UT ST RA Others
Number of Scenarios 200000 200000 200000 17050 200000 105756 160998
NR Score 98.36 93.28 93.60 85.20 91.28 89.33 90.10
@@
Two key insights can be derived from Table 4. Firstly, the quantity of scenarios in the training dataset positively correlates with the expert model’s performance and resulting scores. For instance, scenarios like u-turns and roundabouts, which have fewer instances, exhibit lower scores. Secondly, the complexity of the scene rules affects the model’s learning efficacy, simpler rules, such as those for going straight at an intersection, facilitate better performance. This observation aligns with our expectations. Using scenario-based simulations allows us to identify underperforming scenario types, providing a basis for targeted data enhancement or model optimization in future research.
4.3 Ablation Studies
@@
Table 5: Ablation experiments.
@@
Ablation experiments reveal that interaction-oriented loss and prediction significantly enhance performance, underscoring the importance of interaction. EMOE and scene-specific query also play crucial roles in boosting model performance and these two are used in conjunction. The scene-specific query guides the model’s focus to key scene areas, and EMOE processes that area’s information with a scene-adapted expert. Additionally, Table 5 shows that factoring in interactions yields better predictions for surrounding agents.
5 Conculsion
This paper introduces the EMOE-Planner, a novel framework tailored for enhancing trajectory planning in urban autonomous driving by utilizing scene information and interaction dynamics. The framework employs scene-specific queries generated through multiple scene anchors, focusing the model’s attention on relevant areas. Integration of a Mixture of Experts (MoE) with a shared scene router allows for precise matching of scene types with respective experts, improving the model’s responsiveness to varied scenarios and increasing the transparency of expert selection. Additionally, incorporating interactions between the ego vehicle and other agents into the loss function significantly improves trajectory scoring. Benchmark evaluations on several NuPlan datasets demonstrate that the EMOE-Planner surpasses state-of-the-art (SOTA) models by three points in closed-loop evaluations and closely aligns with the performance of post-processed SOTA methods, reducing reliance on rule-based post-processing.
Acknowledgments and Disclosure of Funding
Funding
This work was supported by internal research funds from BYD Company Limited (No external funding was received).
Competing interests
All authors are employees of BYD Company Limited. The research represents original work without conflicts of interest.
Appendix A Supplementary Material
Figures 4 through 7 illustrate the closed-loop simulation results for four distinct scenarios. We present snapshots at 0s, 5s, 10s, and 15s for both Pluto and EMOE-Planner. In Figure  4(b), it is clear that Pluto collides with a stationary vehicle during a right turn, while EMOE-Planner navigates a safe trajectory. The trajectory by EMOE-Planner in Figure 4(f) initially aligns more closely with the ground truth, attributable to the incorporation of temporal-weighted regression loss, which prioritizes near-term planning and enhances learning from the expert’s obstacle avoidance maneuvers. Conversely, Pluto prioritizes overall trajectory resemblance to the expert, potentially compromising local accuracy for global alignment, leading to less optimal near-term planning.

### v2

@@ Abstract
Despite over a decade of development, autonomous driving trajectory planning in complex urban environments continues to encounter significant challenges. These challenges include the difficulty in accommodating the multi-modal nature of trajectories, the limitations of single expert model in managing diverse scenarios, and insufficient consideration of environmental interactions. To address these issues, this paper introduces the EMoE-Planner, which incorporates three innovative approaches. Firstly, the Explicit MoE (Mixture of Experts) dynamically selects specialized experts based on scenario-specific information through a shared scene router. Secondly, the planner utilizes scene-specific queries to provide multi-modal priors, directing the model’s focus towards relevant target areas. Lastly, it enhances the prediction model and loss calculation by considering the interactions between the ego vehicle and other agents, thereby significantly boosting planning performance. Comparative experiments were conducted on the Nuplan dataset against the state-of-the-art methods. The simulation results demonstrate that our model consistently outperforms SOTA models across nearly all test scenarios. Our model is the first pure learning model to achieve performance surpassing rule-based algorithms in almost all Nuplan closed-loop simulations.
Index Terms:
Autonomous Driving, Trajectory Planning, Imitation Learning, Mixture of Experts.
I Introduction
@@
  • Propose a novel scene-specific learnable query to enable the model to focus on area that the goal point is located. The query simultaneously provides prior information to the model at both the decision level and the modality level, allowing the model to achieve better training efficiency and performance with only a small number of queries.
@@ III-A Overview
Figure 1 illustrates the framework of the proposed model. The model receives the ego vehicle state information $F_{e}\in R^{D_{e}}$, dynamic object information $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, static object information $F_{s}\in R^{N_{s}\times D_{s}}$, and map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$. These inputs are processed through respective encoding modules to obtain D-dimensional features. The details of these input encoding modules will be elaborated in Section 3.2. It is worth noting that the positional and heading angle information in the input data are relative values defined in the ego vehicle’s coordinate system.
Figure 1: Framework of EMoE-Planner
@@
III-B Feature Encoders
@@
As Equation 3 shows, SDE [5] and multi-layer perceptrons (MLP) are employed to encode $F_{e}$ and $F_{s}$. The ego status $F_{e}\in R^{D_{e}}$, where $D_{e}$ includes positional and heading information, is embedded into dimention $D$. Similarly, static object information $S\in R^{N_{s}\times D_{s}}$, where $N_{s}$ is the number of static objects and $D_{s}$ includes attributes such as position, heading, type, and shape, is also embedded into D for concatenation in later stages. $E_{ego}\in R^{D}$ and $E_{static}\in R^{N_{s}\times D}$ are encoded features of ego staus and static objects.
@@
Unlike EE and ES, the encoding of agent and map information requires consideration of both temporal and spatial compression. Equation 4 and 5 shows the process of EA and EM. Raw input of agent is $F_{a}\in R^{N_{a}\times T_{h}\times D_{a}}$, where $N_{a}$ is the number of agents, $T_{h}$ the historical time steps, and $D_{a}$ encompasses attributes such as position, heading, velocity, shape, and type. The ecoded feature of dynamic agents is $E_{agent}\in R^{N_{a}\times D}$. Dimention $D_{a}$ is first embedded into dimention $D$ with Fourier embeddings (FoPE) [48], followed by a compression of $N_{a}$ to a singular dimension via MLP-Mixer [49]. Fope effectively capture periodic patterns in the data, facilitating effective temporal compression. Subsequently, the MLP-Mixer employs dual layers of MLPs to perform interactions both within and across the temporal dimension. Similarly, the map information $F_{m}\in R^{N_{m}\times L_{p}\times D_{m}}$ is also processed with two steps. But the first step is not with fope but a normal linear layer, which is better to handle spacial information. The ecoded feature of map infomation is $E_{map}\in R^{N_{m}\times D}$.
$E_{agent}=\text{MLP-Mixer}(\text{FoPE}(F_{a}))$ (4)
$E_{map}=\text{MLP-Mixer}(\text{Linear}(F_{m}))$ (5)
III-C Explicit MOE
@@
Our approach deviates from vanilla MoE models by introducing an architecture designed specifically for our task. Figure 2(a) depicts a standard MoE model where at each layer, a router selects top $N_{E}$ best-suited expert from all experts based on the input information, with $N_{E}=1$ in our case, to generate the output. However, this architecture has several limitations: the router lacks explicit constraints, making its selection logic difficult to interpret and prone to confusion in similar scenarios. It also risks an uneven distribution of expert utilization.
@@
Figure 2: Differences between vanilla MoE and EMoE
@@
III-D Scene-specific query
@@
III-E Interaction-oriented loss and prediction
III-E1 Interaction-oriented loss
@@
The method for extracting interaction intervals is as follows: traverse each time step $t$ and assess all trajectory points of other agents for potential collisions based on their shapes and orientation angles. Subsequently, using the results from the collision checks, we determine the beginning and ending of interaction intervals, $t_{in},t_{out}\in[1:T_{f}]$. It is essential to clarify that the collisions mentioned herein do not pertain to actual physical interactions between vehicles but rather to instances where two vehicles occupy the same spatial location at different time steps. Furthermore, interaction labels (overtake/yield) are assigned based on the temporal sequence in which the vehicles arrive at the location. Specifically, if the ego vehicle at time $t_{e}$ “collides” with the i-th vehicle at time $t_{a}$, and $t_{e}<t_{a}$ (i.e., the ego vehicle passes through first), then $t_{in}=t_{e}$, and the interacting vehicle is labeled as ”overtake”.
For the i-th agent, the process of $t_{\text{in}}^{i}$ and $t_{\text{out}}^{i}$ is shown in Equation 8-9. The interaction interval is defined as the union of all ($t_{\text{in}}^{i}$,$t_{\text{out}}^{i}$) intervals, as shown in Equation 10.
$\begin{split}t_{\text{in}}^{i}=\min\bigl{\{}t_{a,i}&\mid\exists t_{e}\in\{1,\ldots,T_{f}\},\\ &\quad\|y_{a,i}^{t_{a,i}}-y_{e}^{t_{e}}\|<d_{\text{thr}}\bigr{\}}\end{split}$ (8)
$\begin{split}t_{\text{out}}^{i}=\max\bigl{\{}t_{a,i}&\mid\exists t_{e}\in\{1,\ldots,T_{f}\},\\ &\quad\|y_{a,i}^{t_{a,i}}-y_{e}^{t_{e}}\|<d_{\text{thr}}\bigr{\}}\end{split}$ (9)
@@
Based on the temporal-weighted regression loss in Equation 7, we set the weights of the ego trajectory points within the interaction intervals to 1 to highlight their importance.
@@
III-E2 Interaction-oriented prediciton
@@
IV Experiments
IV-A Experimental Setup
@@
We trained all models on eight Nvidia H800 graphics cards, each with 80GB of VRAM, setting the batch size to 32 and training for 35 epochs. The training time for one million data is approximately 16 hours. The parameters used for training are listed in Table I.
TABLE I: Parameters Used in EMOE-Planner
Name Description Values
lr learning rate 1e-3
$T_{f}$ Future Steps 80
$T_{h}$ History Steps 20
$L_{e}$ Num. scene encoder layers 4
$L_{d}$ Num. decoder layers 4
$N_{E}$ Num. experts 7
$K_{a}$ Num. anchors 24
$k_{R}$ Weight decay factor 0.2
$D$ Dimention of features 128
$D_{h}$ Hidden dim of expert 512
IV-B Results
We assess the performance of all baseline algorithms and EMOE-Planner in NR simulations across Val14, Test14, and Test14-hard. Table II presents the evaluation outcomes while Table III details the metric results for the top two baseline methods and our method.
TABLE II: Closed-loop Results. Higher scores indicate better performance.
@@ IV-B Results
TABLE III: Metrics Results.
@@
According to the results in Table II, our model outperforms all other learning-based methods across all testing scenarios, particularly excelling in the val14 and test14-hard scenarios where it exceeds the state-of-the-art (SOTA) results by a minimum of 3 points. In the val14 scenarios, the performance of the EMOE-Planner closely approaches that of expert scores.Among rule-based methods, when implementing the same post-processing technique used in Pluto and DiffusionPlanner with our EMOE-Planner, it achieves the highest performance in the val14 and test14-hard scenarios, albeit slightly lower than DiffusionPlanner in the test14 scenario. It is noteworthy that the performance of EMOE-Planner with and without post-processing shows slight differences, suggesting that the trajectories produced by the model are of high quality, potentially diminishing the necessity for post-processing, aligning with our objectives. Additionally, the EMoE-Planner, without post-processing, scores only 0.2 points lower than PDM-Closed, which is a rule-based but state-of-the-art algorithm in the NuPlan Challenge, on the val14 test set. On all other datasets, it outperforms PDM-Closed. To our knowledge, we are the first pure learning model to achieve superior performance in closed-loop simulations compared to rule-based algorithms on the NuPlan dataset. In comparison, only when post-processing is applied do the Pluto and Diffusion Planners achieve scores higher than those of PDM-Closed. Furthermore, on the Test14-hard dataset, EMOE-Planner significantly outperforms PDM-Closed. This indicates that rule-based algorithms struggle to handle complex scenarios, while our model continues to perform well in these situations.
@@
As indicated in Table III, our model generally outperforms both Pluto and DiffusionPlanner in all metrics except speed. Notably, it excels in comfort and progress metrics. The low progress score of Pluto aligns with observations from simulations where Pluto’s trajectories tend to be conservative, often slower than expert trajectories. In the appendix, we provide detailed trajectories for various scenarios to give a comprehensive understanding of the situations discussed here. Our loss function accounts for vehicle interactions, which effectively optimizes the trajectory progress, particularly in overtaking scenarios, thus avoiding excessive conservatism while preventing collisions.
@@
In the EMOE architecture, the distribution of scenario quantities in the training data is crucial for model performance. The second row in Table IV details the scenario counts in the training dataset. ST-J, LT-J, RT-J means go straight, turn left and turn right in junctions. ST means go straight in non-junction areas. UT is u-turn and RA is roundabout. During training, we capped each scenario type at 200,000 from a pool of 3,000,000 samples as discussed in section 4.1. The first four scenario types reached this cap, while the last three did not due to insufficient data in the 3M training database, resulting in a total of 1,083,804 samples used for training. We categorized the scenarios in val14 using the method outlined in the appendix and conducted separate simulation assessments for each scenario type, with results presented in Table IV.
TABLE IV: Results of different scenarios.
Scenario Type Number Score Scenario Type Number Score
ST-J 200000 98.36 LT-J 200000 93.28
RT-J 200000 93.60 UT 17050 85.20
ST 200000 91.28 RA 105756 89.33
Others 160998 90.10 NA NA NA
@@
Two key insights can be derived from Table IV. Firstly, the quantity of scenarios in the training dataset positively correlates with the expert model’s performance and resulting scores. For instance, scenarios like u-turns and roundabouts, which have fewer instances, exhibit lower scores. Secondly, the complexity of the scene rules affects the model’s learning efficacy, simpler rules, such as those for going straight at an intersection, facilitate better performance. This observation aligns with our expectations. Using scenario-based simulations allows us to identify underperforming scenario types, providing a basis for targeted data enhancement or model optimization in future research.
IV-C Ablation Studies
@@
TABLE V: Ablation experiments.
@@
Ablation experiments reveal that interaction-oriented loss and prediction significantly enhance performance, underscoring the importance of interaction. EMOE and scene-specific query also play crucial roles in boosting model performance and these two are used in conjunction. The scene-specific query guides the model’s focus to key scene areas, and EMOE processes that area’s information with a scene-adapted expert. Additionally, Table V shows that factoring in interactions yields better predictions for surrounding agents.
IV-D Simulation Results
Figures 4 through 7 illustrate the closed-loop simulation results for four distinct scenarios. We present snapshots at 0s, 5s, 10s, and 15s for both Pluto and EMOE-Planner. In Figure  4(b), it is clear that Pluto collides with a stationary vehicle during a right turn, while EMOE-Planner navigates a safe trajectory. The trajectory by EMOE-Planner in Figure 4(f) initially aligns more closely with the ground truth, attributable to the incorporation of temporal-weighted regression loss which prioritizes near-term planning and enhances learning from the expert’s obstacle avoidance maneuvers. Conversely, Pluto prioritizes overall trajectory resemblance to the expert, potentially compromising local accuracy for global alignment, leading to less optimal near-term planning.